Controls Meeting 17-May-2012

From GlueXWiki
Jump to: navigation, search


  • Announcements
    • CNU summer students
  • Minutes from previous meeting 3-May-2012
  • Barcode scanner - Elliott, Hovanes
  • FCC Class B UPS's - Hovanes, Elliott
  • EPICS - Hovanes, Vanik
    • alarm system, goniometer, UConn boards, other power supplies, etc.
  • Counting house installation - Hovanes, Mark
  • Solenoid_Controls_Redesign - Mark, Josh, Scot, Yi, Elliott
    • FMEA
    • UPS power to Danfysik, SOE, QD, overcurrent protection, ground fault detector, etc.
  • JInventory database - Elliott
  • Test stand - Hovanes, Josh
  • PLC programming and PXI system - Josh
  • Elogs and databases - Elliott
  • Mantis task tracker
  • Experiment Control Supervisor - Elliott


9:30 AM Thurs 17-May-2012 CC L207


Next Meeting

9:30am 31-May-2012 CC L207


Present: Elliott W, Hovanes E, Vanik K, Mark S, Dave B.

JInventory system and barcode scanner

  • Hovanes suggested adding two new fields, MAC_Address and IPName. Elliott will do this. The Telecom field will stay.
  • Wired barcodes scanners around $300, full-function wireless around $1600 or more.
  • Elliott will find out about Bluetooth versions.
  • Consensus is that we need at least two, one wired, one wireless.
  • Will purchase wired model now, not sure whether to buy wireless model now or later.
  • Consensus to enter slots as full items in the JInventory system.
  • Elliott will create scripts to automate creation of slot entries for crates.
  • Josh entered many A-B modules.
  • Mark will enter electrical substations and panel boards.

Small UPS's

  • Still not clear how much power needed. We know what out equipment needs, not sure about other systems (e.g. SiPM cooling).
  • Hovanes found high-capacity expensive FCC Class B UPS's at Grainger. Low-capacity ones are easy to find.
  • Elliott asked Paul Powers to look into filtering Class A power down to Class B specs.
  • We believe our power factors are typically near one, maybe 0.9.


  • Vanik continues to make progress on the stepper motor controller. He got the optical counter feature to work.
  • Will likely need optical limit switches since "home" feature not working well. Unfortunately they are easily radiation damaged.
  • UConn boards may be redesigned.
  • We're hoping Accelerator will take over four of our five power supplies (not the Solenoid supply).

Counting House

  • RDC will be installed next week.
  • Giles finalizing rack room power distribution plan.
  • Everything is cleaned up for the Open House.


  • Mark installed new Danfysik boards.
  • Mark installed center tap on dump resistor.
  • Mark is working on SOE circuits.
  • We misplaced Hall C's flow switch, had to order another.

Test Stand

  • Hovanes ordered a 32-port terminal server, updated version of the ones used in Hall B. About $1500.


  • Dave Butler is learning about our PLC system.
  • Josh may have a good first project for him, reworking the Danfysik communication code.